# Towards a test of the Born rule in high-energy collisions

### v2

@@
Theoretical Physics, Blackett Laboratory, Imperial College London,
@@
We consider how the Born rule, a fundamental principle of quantum mechanics, could be tested for particles created on the shortest timescales ($\sim 10^{-25}\,\mathrm{s}$) currently accessible at high-energy colliders. We focus on targeted tests of the Born rule for spin or polarisation probabilities, which offer a particularly clean experimental signal, and which can be described by a simple hidden-variables model of two-state systems proposed by Bell. These probabilities test a remarkable feature of the quantum formalism, whereby expectation values for incompatible experiments are linearly related. Born-rule violations can be parameterised by nonlinear expectation values for quantum measurements of spin or polarisation, along with anomalies in ensemble averages, which are to be constrained by experiment. Notable experiments considered here include the recent detection of single photons from top-quark decay, and the indirect measurement of tau-lepton polarisation. Repurposing these experiments as tests of the Born rule, however, presents several challenges, which are discussed in this paper.
@@ 1 Introduction
More recently it has been argued that, in the de Broglie-Bohm formulation of canonical quantum gravity, the Born rule may be gravitationally unstable at the Planck scale [37, 38]. Such effects could in principle manifest today as a small violation of the Born rule in Hawking radiation from exploding primordial black holes, whose gamma-ray emissions are being searched for [39]. For such rays, however, tests of the Born rule would require a dedicated satellite in space to avoid absorption by the earth’s atmosphere. This and related suggestions have motivated a test of the Born rule aboard the forthcoming QUICK3 satellite mission, which includes an interferometer designed to test the Born rule for photons from an onboard optical source [40]. Once the proof-of-concept experiment is fully operational, in the future it could be redeployed to test the Born rule for photons from exotic astrophysical sources, such as exploding primordial black holes [37, 38] or disintegrating dark matter [41].
@@
A similar experiment was in fact carried out in 1967 by Papaliolios [48], with optical photons from a tungsten filament lamp. The purpose was to test a hidden-variables model due to Bohm and Bub [49] (distinct from de Broglie-Bohm theory), in which deviations from the Born rule are predicted to exist for a very short time (perhaps of order $10^{-13}\,\mathrm{s}$) immediately after a quantum measurement has taken place. According to the model, if a second quantum measurement takes place within (say) $\sim 10^{-13}\,\mathrm{s}$ of the first, deviations from the Born rule should be observed. To test this, Papaliolios placed three linear polarisers in close proximity and varied the angle of the third. The first polariser prepared a linearly-polarised quantum state, the second measured the polarisation along an axis almost orthogonal to the first, while the third performed a second measurement at an angle $\Theta$ relative to the first measurement. By placing the polarisers sufficiently close together to guarantee the short timescale, Papaliolios verified that the transmission probability through the third polariser varied with $\Theta$ according to (3) within an accuracy of about $1\%$, thereby confirming the approximate validity of the Born rule over timescales of order $10^{-13}\,\mathrm{s}$ [^5] Historically, this appears to have been the first experiment designed to test a hidden-variables theory [50].
@@
For completeness, we briefly summarise the state-of-the-art in gamma-ray polarimetry. The widely-used technique of Compton polarimetry exploits the fact that, in Compton scattering, photons are scattered preferentially perpendicular to the incident mean polarisation vector, a technique that works well in the energy range up to about 100 KeV [52]. Recent progress in gamma-ray polarimetry is exemplified by upcoming instruments such as POLAR-2 [53], which will use Compton scattering in the soft gamma-ray region to extract mean polarisation from astrophysical gamma-ray bursts, by analysing the distribution of azimuthal scattering angles over many events. Proposed missions, including AMEGO, AMEGO-X, and e-ASTROGAM [54, 55, 56] aim to measure the mean polarisation of gamma rays in the MeV-GeV range by combining Compton scattering with electron–positron pair production. However, while all three proposed missions incorporate the necessary instrumentation to study polarisation across this energy range, their projected polarimetric sensitivity remains limited to the MeV regime, where Compton-based methods remain the most effective. It may be hoped that, in the near future, similar techniques might be applied to the study of polarisation for gamma rays emerging from high-energy collisions.
In Section 2, we discuss the Born rule for two-state quantum systems. We outline the formalism of hidden-variables theories and show how they imply an extended physics with Born-rule violations parameterised by anomalous expectation values. These ideas are illustrated by the simple Bell model of two-state systems. In Section 3, we discuss the detection of photons emerging from short-timescale collider processes, in particular single photons from top-quark decay and photon pairs from neutral pion decay. We consider how such experiments might be adapted to test the Born rule for photon polarisation probabilities. Section 4 focuses on average polarisation measurements for tau leptons from $Z$-boson decay and considers how these could be used to study possible deviations from the Born rule. In Section 5, we highlight how collider experiments employ the Born rule for background modelling and event selection, and we discuss the implications this may have for our proposed tests. Finally, in Section 6, we summarise our findings and discuss the broader significance of possible Born-rule violations at high energies.
@@ 2 The Born rule for two-state systems
For particles of spin $\tfrac{1}{2}$, $\sigma=\pm 1$ indicates spin (in units /2Planck-constant-over-$\hbar/2$) up or down a spatial axis $\mathbf{m}$. For photons, $\sigma=\pm 1$ indicates polarisation parallel or perpendicular to a spatial axis $\mathbf{M}$, where now $\theta$ corresponds to an angle $\Theta=\theta/2$ in 3-space.
@@ 2.1 Hidden variables and the Born rule
[…]
$\left\langle\omega\right\rangle_{\mathrm{QT}}=\mathrm{Tr}(\hat{\rho}\hat{\omega})$ (8)
@@
[…]
@@
[…]
@@ 2.2 Parameterising deviations from the Born rule
with $\mathbf{m}$ specified by angular coordinates $(\theta,\phi)$ on the Bloch sphere and $\mathbf{P}_{\mathrm{QT}}$ pointing along $+z$ [^7] The reality of $p^{+}$ implies $b_{lm}^{\ast}=(-1)^{m}b_{l(-m)}$. In quantum theory, $p^{+}=p_{\mathrm{QT}}^{+}$ takes the form (6) with a simple dipole term in $\mathbf{m}$, and the only non-zero coefficients are
$(b_{00})_{\mathrm{QT}}=\sqrt{\pi}\ ,\ (b_{10})_{\mathrm{QT}}=P\sqrt{\pi/3}\ .$ (13)
@@ 2.3 The Bell model
[…]
@@
[…]
@@
[…]
@@
[…]
@@
(again with the $z$-axis defined by $\mathbf{P}_{\mathrm{QT}}$). For photon polarisation measurements along an angle $\Theta=\theta/2$ in 3-space, the usual Malus law (3) tested by Papaliolios then receives a correction
$\Delta p^{+}(\Theta)=\frac{1}{2}R_{zzz}\cos^{3}(2\Theta)\,,$ (27)
@@ 3.1 Top-quark decay
Recent measurements by the ATLAS Collaboration [59] have successfully demonstrated this technique, detecting $5{,}734$ photons from top-quark events with a total integrated luminosity of $139\,\mathrm{fb}^{-1}$. The experiment distinguished between photons radiated directly from the top quark ($4{,}899$ events) and those from decay products ($835$ events) using the mass-reconstruction method described above.
To implement this approach, the ATLAS analysis relies on Monte Carlo (MC) simulations to estimate both the signal and background processes. A neural network (NN) is then trained on these MC-generated events, learning how to distinguish photons radiated directly by the top quark from those originating in the decay products. This trained NN is subsequently applied to real data, using reconstructed kinematics (including mass-related observables) to classify whether a photon most likely comes from the top quark itself or from a decay product.
We propose to build upon this successful measurement by adding polarisation measurements, whether for individual photons or only ensemble-averages. These measurements must be non-invasive, ensuring that all other parts of the ATLAS detector remain unchanged. The ATLAS detector [60] identifies photons by combining information from multiple subsystems (Fig. 4). First, the inner tracking detector ensures there is no charged-particle track associated with the candidate photon. Next, the electromagnetic (EM) calorimeter measures a localised electromagnetic shower. Finally, an isolation requirement is imposed: within a small angular cone around the photon candidate, the total energy measured in both the EM and hadronic calorimeters must remain below a specified threshold. This helps distinguish single photons from more energetic hadronic jets.
@@ 3.2 Neutral pion decay
An alternative advantageous example for testing the Born rule is neutral pion decay. The neutral pion, with a mass of approximately 135 MeV [58], decays into two photons with a lifetime of about $8.4\times 10^{-17}\,\mathrm{s}$. Although this timescale is much longer than that of the top quark, it remains four orders of magnitude shorter than the timescale probed in the Papaliolios experiment [48]. The $\pi^{0}$ system offers a crucial advantage: it decays almost exclusively into two photons (with a branching ratio of about $98.8\%$), minimising ambiguities in particle identification such as those associated with top-quark decay.
Typically, $\pi^{0}$ mesons are produced in proton-proton ($pp$) collisions [61]. Since $pp$ collisions produce a variety of particles, one must separate the photons originating from $\pi^{0}$ decay ($\pi^{0}\rightarrow\gamma\gamma$) from those arising via other processes including accidental coincidences between photons from different $\pi^{0}$’s, direct photons, photons from $\eta$ decay, and misidentified neutrons. The procedure is analogous to that used for top quarks: if both photons originate from the same $\pi^{0}$, their invariant mass will reconstruct to $135\,\mathrm{MeV}$. If the two-photon mass falls outside a window of approximately ±$\pm 3\sigma$ around this peak, the photon pair is classified as background.

### v4

@@
Abdus Salam Centre for Theoretical Physics, Imperial College London,
@@
We consider how the Born rule, a fundamental principle of quantum mechanics, can be tested for particles created on the shortest timescales ($\sim 10^{-25}\,\mathrm{s}$) currently accessible at high-energy colliders. We focus on targeted tests of the Born rule for spin or polarisation probabilities, which offer a particularly clean experimental signal, and which can be described by a simple hidden-variables model of two-state systems proposed by Bell. These probabilities test a remarkable feature of the quantum formalism, whereby expectation values for incompatible experiments are linearly related. Born-rule violations can be parameterised by nonlinear expectation values for quantum measurements of spin or polarisation, together with anomalies in ensemble averages, which may then be constrained by experiment. Notable experiments considered here include the recent detection of single photons from top-quark decay, and the indirect measurement of tau-lepton polarisation. Repurposing these experiments as tests of the Born rule, however, presents several challenges, which are discussed in this paper.
@@ 1 Introduction
More recently it has been argued that, in the de Broglie-Bohm formulation of canonical quantum gravity, the Born rule may be gravitationally unstable at the Planck scale [37, 38]. Such effects could in principle manifest today as a small violation of the Born rule in Hawking radiation from exploding primordial black holes, whose gamma-ray emissions are being searched for [39]. For such rays, however, tests of the Born rule would require a dedicated satellite in space to avoid absorption by the earth’s atmosphere. This and related suggestions have motivated a test of the Born rule aboard the recently launched QUICK3 satellite mission, which includes an interferometer designed to test the Born rule for photons from an onboard optical source [40]. Once the proof-of-concept experiment is fully operational, in the future it could be redeployed to test the Born rule for photons from exotic astrophysical sources, such as exploding primordial black holes [37, 38] or disintegrating dark matter [41].
@@
A similar experiment was in fact carried out in 1967 by Papaliolios [48], with optical photons from a tungsten filament lamp. The purpose was to test a hidden-variables model due to Bohm and Bub [49] (distinct from de Broglie-Bohm theory), in which deviations from the Born rule are predicted to exist for a very short time (perhaps of order $10^{-13}\,\mathrm{s}$) immediately after a quantum measurement has taken place. According to the model, if a second quantum measurement takes place within (say) $\sim 10^{-13}\,\mathrm{s}$ of the first, deviations from the Born rule should be observed. To test this, Papaliolios placed three linear polarisers in close proximity and varied the angle of the third. The first polariser prepared a linearly-polarised quantum state, the second measured the polarisation along an axis almost orthogonal to the first, while the third performed a second measurement at an angle $\Theta$ relative to the first measurement. By placing the polarisers sufficiently close together to guarantee the short timescale, Papaliolios verified that the transmission probability through the third polariser varied with $\Theta$ according to (3) within an accuracy of about $1\%$, thereby confirming the approximate validity of the Born rule over timescales of order $10^{-13}\,\mathrm{s}$. [^5] Historically, this appears to have been the first experiment designed to test a hidden-variables theory [50].
@@
For completeness, we briefly summarise the state-of-the-art in gamma-ray polarimetry. The widely used technique of Compton polarimetry exploits the fact that, in Compton scattering, photons are scattered preferentially perpendicular to the incident mean polarisation vector, a technique that works well in the energy range up to about 100 keV [52]. Recent progress in gamma-ray polarimetry is exemplified by upcoming instruments such as POLAR-2 [53], which will use Compton scattering in the soft gamma-ray region to extract mean polarisation from astrophysical gamma-ray bursts, by analysing the distribution of azimuthal scattering angles over many events. Proposed missions, including AMEGO, AMEGO-X, and e-ASTROGAM [54, 55, 56] aim to measure the mean polarisation of gamma rays in the MeV-GeV range by combining Compton scattering with electron–positron pair production. However, while all three proposed missions incorporate the necessary instrumentation to study polarisation across this energy range, their projected polarimetric sensitivity remains limited to the MeV regime, where Compton-based methods remain the most effective. It may be hoped that, in the near future, similar techniques might be applied to the study of polarisation for gamma rays emerging from high-energy collisions.
In Section 2, we discuss the Born rule for two-state quantum systems. We outline the formalism of hidden-variables theories and show how they imply an extended physics with Born-rule violations parameterised by anomalous expectation values. These ideas are illustrated by the simple Bell model of two-state systems. In Section 3, we discuss the detection of photons emerging from short-timescale collider processes, in particular single photons from top-quark decay and photon pairs from neutral pion decay. We consider how such experiments might be adapted to test the Born rule for photon polarisation probabilities. Section 4 focuses on average polarisation measurements for tau leptons from $Z$-boson decay and considers how these may be used to study possible deviations from the Born rule. In Section 5, we highlight how collider experiments employ the Born rule for background modelling and event selection, and we discuss the implications this may have for our proposed tests. Finally, in Section 6, we summarise our findings and discuss the broader significance of possible Born-rule violations at high energies.
@@ 2 The Born rule for two-state systems
For particles of spin $\tfrac{1}{2}$, $\sigma=\pm 1$ indicates spin (in units $\hbar/2$) up or down a spatial axis $\mathbf{m}$. For photons, $\sigma=\pm 1$ indicates polarisation parallel or perpendicular to a spatial axis $\mathbf{M}$, where now $\theta$ corresponds to an angle $\Theta=\theta/2$ in 3-space.
@@ 2.1 Hidden variables and the Born rule
A deterministic hidden-variables theory posits unknown parameters $\lambda$ that determine the outcomes of quantum measurements. Formally, for an experiment specified by apparatus settings $M$, such a theory provides a mapping $\omega=\omega(M,\lambda)$ from initial values of $\lambda$ to final outcomes $\omega$ (equal to eigenvalues of the operator observable $\hat{\omega}$ that is being measured). Over an ensemble of experiments with fixed $M$, the hidden variables $\lambda$ will generally vary from one run of the experiment to the next. For a general ensemble distribution $\rho(\lambda)$, the expectation value of $\omega$ will be [^6] Here $\int d\lambda$ is understood as a generalised sum, in case $\lambda$ is not continuous.
$\left\langle\omega\right\rangle=\int d\lambda\ \rho(\lambda)\omega(M,\lambda)\ .$ (7)
@@
$E_{\mathrm{QT}}(\mathbf{m})\equiv\left\langle\sigma\left(\mathbf{m},\lambda\right)\right\rangle_{\mathrm{QT}}=\int d\lambda\ \rho_{\mathrm{QT}}(\lambda)\sigma\left(\mathbf{m},\lambda\right)=\mathbf{m}\cdot\mathbf{P}_{\mathrm{QT}}\,.$ (10)
@@
$E(\mathbf{m})\equiv\left\langle\sigma\left(\mathbf{m},\lambda\right)\right\rangle=\int d\lambda\ \rho(\lambda)\sigma\left(\mathbf{m},\lambda\right)$
@@ 2.2 Parameterising deviations from the Born rule
with $\mathbf{m}$ specified by angular coordinates $(\theta,\phi)$ on the Bloch sphere and $\mathbf{P}_{\mathrm{QT}}$ pointing along +z+z. [^7] The reality of $p^{+}$ implies $b_{lm}^{\ast}=(-1)^{m}b_{l(-m)}$. In quantum theory, $p^{+}=p_{\mathrm{QT}}^{+}$ takes the form (6) with a simple dipole term in $\mathbf{m}$, and the only non-zero coefficients are
$(b_{00})_{\mathrm{QT}}=\sqrt{\pi}\ ,\ (b_{10})_{\mathrm{QT}}=P_{\mathrm{QT}}\sqrt{\pi/3}\ .$ (13)
@@ 2.3 The Bell model
$E_{\mathrm{QT}}(\mathbf{m})=\int d\Omega\ \rho_{\mathrm{QT}}(\boldsymbol{\lambda})\sigma\left(\mathbf{m},\boldsymbol{\lambda}\right)=\mathbf{m}\cdot\mathbf{P}_{\mathrm{QT}}=\cos\theta$ (18)
@@
$p_{\mathrm{QT}}^{-}(\mathbf{m})=\int_{A_{-}}d\Omega\ \rho_{\mathrm{QT}}(\boldsymbol{\lambda})=\frac{1}{4\pi}A_{-}\,$, (19)
@@
$p_{\mathrm{QT}}^{+}(\mathbf{m})=\int_{A_{+}}d\Omega\ \rho_{\mathrm{QT}}(\boldsymbol{\lambda})=\frac{1}{4\pi}A_{+}\,$,
@@
$E(\mathbf{m})=p^{+}(\mathbf{m})-p^{-}(\mathbf{m})=\int d\Omega\ \rho(\boldsymbol{\lambda})\sigma\left(\mathbf{m},\boldsymbol{\lambda}\right)$ (23)
@@
(again with the $z$-axis defined by $\mathbf{P}_{\mathrm{QT}}$). However, from (20), for $\theta=0$ we have $A_{-}=0$ while for $\theta=\pi$ we have $A_{+}=0$. The total transmission probability $p^{+}(\theta)$ is therefore subject to the constraints $p^{+}(0)=1$ and $p^{+}(\pi)=0$ for all finite distributions $\rho$. To satisfy these constraints, we must include a correction to the effective mean polarisation as well, with an anomalous expectation value
$E(\theta)=(1-R_{zzz})\cos\theta+R_{zzz}\cos^{3}\theta\,.$ (27)
This corresponds to a corrected transmission probability
$p^{+}(\theta)=p_{\mathrm{QT}}^{+}(\theta)-\frac{1}{2}R_{zzz}\cos\theta+\frac{1}{2}R_{zzz}\cos^{3}\theta\,.$ (28)
Note that the symmetry $E(-\mathbf{m})=-E(\mathbf{m})$ requires $E=0$ for $\theta=\pi/2$, hence $p^{+}(\pi/2)=\frac{1}{2}$, which is consistent with (28).
For photon polarisation measurements along an angle $\Theta=\theta/2$ in 3-space, we may write the corrected transmission probability in the form
$p^{+}(\Theta)=p_{\mathrm{QT}}^{+}(\Theta)+\frac{1}{2}R_{zzz}f(\Theta)\,$, (29)
with
$f(\Theta)=-\cos(2\Theta)\sin^{2}(2\Theta)$ (30)
(where $f(\Theta)=0$ at $\Theta=0,\pi/4,\pi/2$). The usual Malus law (3) tested by Papaliolios then receives a correction
$\Delta p^{+}(\Theta)=\frac{1}{2}R_{zzz}f(\Theta)\,$, (31)
@@ 3.1 Top-quark decay
Recent measurements by the ATLAS Collaboration [59] have successfully demonstrated this technique, corresponding to a signal of 5,7345{,}734 photons from top-quark events with a total integrated luminosity of $139\,\mathrm{fb}^{-1}$, as inferred from the associated HEPData tables. The experiment distinguished between photons radiated directly from the top quark (4,8994{,}899 events) and those from decay products ($835$ events) using the mass-reconstruction method described above.
To implement this approach, the ATLAS analysis relies on Monte Carlo (MC) simulations to estimate both the signal and background processes. A neural network (NN) is then trained on these MC-generated events, learning how to distinguish events with photons from single-top-quark production (where the photon is radiated either directly by the top quark or by a top decay product) from other background processes. This trained NN is subsequently applied to real data, using reconstructed kinematics (including mass-related observables) to classify events as signal-like or background-like.
We propose to build upon this successful measurement by adding polarisation measurements, whether for individual photons or only ensemble averages. These measurements must be non-invasive, ensuring that all other parts of the ATLAS detector remain unchanged. The ATLAS detector [60] identifies photons by combining information from multiple subsystems (Fig. 4). First, the inner tracking detector ensures there is no charged-particle track associated with the candidate photon. Next, the electromagnetic (EM) calorimeter measures a localised electromagnetic shower. Finally, an isolation requirement is imposed: within a small angular cone around the photon candidate, the total energy measured in both the EM and hadronic calorimeters must remain below a specified threshold. This helps distinguish single photons from more energetic hadronic jets.
@@ 3.2 Neutral pion decay
Another useful example for testing the Born rule is neutral pion decay. The neutral pion, with a mass of approximately 135 MeV [58], decays into two photons with a lifetime of about $8.4\times 10^{-17}\,\mathrm{s}$ (Fig. 5). Although this timescale is much longer than that of the top quark, it remains four orders of magnitude shorter than the timescale probed in the Papaliolios experiment [48]. The $\pi^{0}$ system offers a crucial advantage: it decays almost exclusively into two photons (with a branching ratio of about $98.8\%$), minimising ambiguities in particle identification such as those associated with top-quark decay.
Typically, $\pi^{0}$ mesons are produced in proton-proton ($pp$) collisions [61]. Since $pp$ collisions produce a variety of particles, one must separate the photons originating from $\pi^{0}$ decay ($\pi^{0}\rightarrow\gamma\gamma$) from those arising via other processes including accidental coincidences between photons from different $\pi^{0}$’s, direct photons, photons from $\eta$ decay, and misidentified neutrons. The procedure is analogous to that used for top quarks: if both photons originate from the same $\pi^{0}$, their invariant mass will reconstruct to $135\,\mathrm{MeV}$. If the two-photon mass falls outside a window of approximately $\pm 3\sigma$ around this peak, the photon pair is classified as background.
@@
3.3 Experimental constraints and targets
For mean polarisation measurements, for which there already exists technology widely used in astrophysics (Section 1), we might consider a small shift in the magnitude of the mean polarisation,
$P=P_{\mathrm{QT}}+\varepsilon,$ (33)
where $P$, $P_{\mathrm{QT}}$ are respectively the magnitudes of $\mathbf{P}$, $\mathbf{P}_{\mathrm{QT}}$ and $\varepsilon$ is a small parameter. As noted in Section 1, however, current techniques allow us to measure mean photon polarisations at energies rather lower than those required at a collider. As will be discussed in Section 4.4, mean polarisations have in fact already been measured for tau leptons, and for that case current experiments already allow us to put simple bounds on possible deviations from the quantum mean. Here, for the case of photons, we focus on a more ambitious long-term possibility, in which measurements of individual photon polarisations would allow a full probe of the Born rule.
We now consider how accurately individual photon polarisations must be measured in order to place quantitative constraints on the Born rule. We provide bounds for: (i) the minimum number of events $N_{\min}$, (ii) the required transmission count accuracy $\Delta n$, and (iii) the allowed polariser angle misalignment $\Delta\Theta$. Our results provide what seem reasonable targets for the future development of single-photon gamma-ray polarimetry.
From Section 2.3, we know that for photon polarisation measurements along an angle $\Theta=\theta/2$, Malus’ law ($p_{\mathrm{QT}}^{+}(\Theta)=\cos^{2}\Theta$) potentially can receive a correction, which in its simplest form is given by (31). Dropping the subscript zzzzzz, we may write the model probability as
$p^{+}(\Theta;R)=\cos^{2}\Theta+\tfrac{1}{2}Rf(\Theta),$ (34)
where $f(\Theta)$ is given by (30). We wish to consider how to put quantitative experimental constraints on the anomalous parameter $R$, where of course in quantum theory $R=0$.
At a fixed angular setting $\Theta_{\rm{set}}$, we consider the unbiased estimator for $R$,
$\hat{R}(\Theta_{\rm set})\;=\;\frac{2\big(\hat{p}^{+}-\cos^{2}\Theta_{\rm set}\big)}{f\big(\Theta_{\rm set}\big)},$ (35)
with
$\hat{p}^{+}=\frac{n}{N},$ (36)
where $n$ is the number of transmitted photons out of $N$ trials.
3.3.1 Minimum number of events $N_{\min}$
For fixed $\Theta_{\rm{set}}$, $\hat{p}^{+}$ is a binomial random variable with variance
$\sigma_{\hat{p}^{+}}^{2}=\frac{p^{+}(1-p^{+})}{N}.$ (37)
By error propagation, the standard deviation for $\hat{R}$ will be
$\sigma_{\hat{R}}=\Big|\frac{d\hat{R}}{d\hat{p}^{+}}\Big|\sigma_{\hat{p}^{+}}=\frac{2}{|f(\Theta_{\rm set})|}\sqrt{\frac{p^{+}(1-p^{+})}{N}}.$ (38)
In the large-$N$ (Gaussian) limit, the statistical significance associated with an observed value $\hat{R}$ is given by [63]
$Z_{0}=\frac{\hat{R}}{{\sigma_{\hat{R}}}}$. (39)
To claim a $3\sigma$ detection of a deviation $|R|=\delta$, so that $\sigma_{\hat{R}}\leq\delta/3$, we then require
$N\geq\frac{36p^{+}(1-p^{+})}{\delta^{2}{[f(\Theta_{\rm set})}]^{2}}.$ (40)
Taking $p^{+}$ as (to a good approximation) the expected probability under the null hypothesis ($R=0$, so $p^{+}=\cos^{2}\Theta$) gives
$N\geq\frac{36}{\delta^{2}\sin^{2}(4\Theta_{\rm set})}.$ (41)
At $\Theta_{\rm set}=\pi/8$, the right-hand side reaches its minimum so at other angles the minimum number of events will be larger. Taking, for illustration, $\delta=10^{-2}$, this gives
$N_{\min}\geq 3.6\times 10^{5},$ (42)
where $N_{\min}$ refers to the number of signal photons in the absence of background events.
If instead the sample contains a mixture of signal ($N_{s}$) and background ($N_{b}$) events, we have a signal fraction $f_{s}=N_{s}/(N_{s}+N_{b})$ and a total transmission probability
$p^{+}_{\rm tot}=f_{s}\,p^{+}_{s}+(1-f_{s})\,p^{+}_{b}$, (43)
where $p^{+}_{s}$ and $p^{+}_{b}$ are respectively transmission probabilities for signal and background photons. If the background photons show no Born-rule deviation $(R_{b}=0)$, then we find
$p^{+}_{\rm tot}=\cos^{2}\Theta_{\rm set}+\tfrac{1}{2}f_{s}R_{s}f(\Theta_{\rm set}),$ (44)
where $R_{s}$ is the anomalous parameter for the signal photons. This corresponds to an effective dilution of $R$, with $R_{\rm eff}=f_{s}R_{s}$. Since in (41) we found $N_{\rm min}\propto 1/\delta^{2}$, the required number of events in the presence of background becomes
$N_{\rm min}^{\prime}=\frac{1}{f_{s}^{2}}N_{\rm min}.$ (45)
For photons radiated by top quarks, the fraction of signal photons originating from the $tq\gamma$ process can be taken to be $f_{s}\simeq 0.8$ [59], since the $t(\to\ell\nu b\gamma)q$ contribution is found to be $\approx 20\%$ of the events in the fiducial region. Using (45), this gives $N_{\rm min}^{\prime}\approx 5.6\times 10^{5}$. In the experiment, about $5\times 10^{3}$ events were reported, which is two orders of magnitude smaller than our target $N_{\rm min}^{\prime}$. For neutral pion decay, the backgrounds are negligible, and about $2\times 10^{5}$ events were reported in the clean Type-I topology. This corresponds to about $4\times 10^{5}$ photons, slightly exceeding our threshold $N_{\rm min}\geq 3.6\times 10^{5}$.
3.3.2 Transmission count uncertainty $\Delta n$
A small shift in the transmission count $n\rightarrow n+\Delta n$ shifts $\hat{p}^{+}$ by
$\Delta\hat{p}^{+}=\frac{\Delta n}{N},$ (46)
which leads to a shift in the estimator
$\Delta\hat{R}_{\rm(n)}=\frac{2\Delta\hat{p}^{+}}{f(\Theta_{\rm set})}.$ (47)
To prevent systematic effects from dominating the total uncertainty, we require
$|\Delta\hat{R}_{\rm(n)}|\leq\frac{\delta}{3},$ (48)
which corresponds to a bound of
$\Delta\hat{p}^{+}\leq\frac{\delta}{6}|f(\Theta_{\rm set})|.$ (49)
Again taking $\Theta_{\rm set}=\pi/8$, we have $|f(\Theta_{\rm set})|=1/2\sqrt{2}$, hence
$\Delta\hat{p}^{+}\leq\frac{\delta}{12\sqrt{2}},$ (50)
which implies
$\Delta n\leq\frac{\delta}{12\sqrt{2}}N.$ (51)
As before, taking $\delta=10^{-2}$ and with $N=N_{\rm min}=3.6\times 10^{5}$, we have $\Delta n\lesssim 212$. This corresponds to a fractional counting precision of
$\frac{\Delta n}{n}\approx 0.07\%$ (52)
(where $n$ is estimated to be $N\cos^{2}(\pi/8)$).
3.3.3 Angular misalignment $\Delta\Theta$
If the polarimeter is misaligned by a small angle $\Delta\Theta$, so that the true angle
$\Theta_{\rm true}=\Theta_{\rm set}+\Delta\Theta,$ (53)
with $|\Delta\Theta|\ll 1$, then the model probability (34) becomes $p^{+}(\Theta_{\rm true};R)$, and our estimator for $R$ will be
$\hat{R}=\frac{2(p^{+}(\Theta_{\rm true};R)-\cos^{2}\Theta_{\rm set})}{f(\Theta_{\rm set})}=\frac{2\big(\cos^{2}\Theta_{\rm true}-\cos^{2}\Theta_{\rm set}\big)}{f(\Theta_{\rm set})}+R\,\frac{f(\Theta_{\rm true})}{f(\Theta_{\rm set})}.$ (54)
Expanding to first order in $\Delta\Theta$, and neglecting the small $R$-dependent term, gives a shift in the estimator
$\Delta\hat{R}_{\rm(angle)}\simeq-2\frac{\sin(2\Theta_{\rm set})}{f(\Theta_{\rm set})}\Delta\Theta.$ (55)
To ensure a statistically-limited measurement, and to claim a $3\sigma$ significance for a deviation $|R|=\delta$, we again require
$|\Delta\hat{R}_{\rm(angle)}|\leq\frac{\delta}{3},$ (56)
and so we have
$|\Delta\Theta|\leq\frac{\delta}{6}\frac{|f(\Theta_{\rm set})|}{|\sin(2\Theta_{\rm set})|}.$ (57)
At $\Theta_{\rm set}=\pi/8$, and again for $\delta=10^{-2}$, we find a maximum allowed angular misalignment
$|\Delta\Theta|\approx 0.00083\,\text{rad}\approx 0.05^{\circ}.$ (58)
3.3.4 Benchmarks from optical polarimetry
Optical photon polarimetry is a highly developed technology, and the accuracy required from our analysis is not unreasonable compared to what has already been demonstrated in quantum optics. A recent experiment using spontaneous parametric down conversion has measured polarisations for $1.7\times 10^{7}$ individual photons [64]. The analysis showed that to keep polarisation bias below $1\%$, the half-wave plate must be aligned within $0.3^{\circ}$ of the ideal setting. This benchmark is only a factor of six less precise than our requirement (58). In the same experiment, the average number of detections per time bin was measured to be $\mu=0.364\pm 0.002$. This corresponds to a fractional counting precision of $\Delta n/n=0.002/0.364\simeq 0.55\%$, about an order of magnitude less stringent than our target fractional counting precision (52).